# The Skolem Problem in rings of positive characteristic

### v2

@@ 1 Introduction and main results
The Skolem Problem in a commutative ring $\mathcal{R}$ is the problem of determining whether a given linear recurrence sequence $(\gamma_{n})_{n\in\mathbb{N}}$ over $\mathcal{R}$ contains a zero term. The Skolem Problem is deeply related to central questions in program verification [ouaknine2015linear, almagor2021deciding], control theory [blondel2000survey, fijalkow2019decidability], dynamical systems [kannan1986polynomial, agrawal2015approximate], and number theory [bell2006generalised, luca2023skolem]. Despite its simple formulation, decidability of the Skolem Problem over $\mathbb{Z}$ (and equivalently, $\mathbb{Q}$) is a major open problem in mathematics and computer science: see [mahler1935arithmetische, lech1953note, Tijdeman1984, blondel2002presence, lipton2022skolem, luca2025large] for a selected list of partial results and recent progress. Indeed, the celebrated Skolem-Mahler-Lech theorem [skolem1934verfahren] shows that the zero set of a linear recurrence sequence over $\mathbb{Z}$ is a union of a finite set and finitely many arithmetic progressions. However, the Skolem-Mahler-Lech theorem is not effective, meaning there is currently no known computable bound on this finite set, and hence the decidability of the Skolem Problem over the ring $\mathbb{Z}$ remains open.
In contrast, decidability of the Skolem Problem in fields of prime characteristic (such as $\mathbb{F}_{p}(X)$) was shown by Derksen [derksen2007skolem], who gave an effective description of the zero set as a $p$-normal set. In this paper we continue Derksen’s work in this direction by showing decidability of the Skolem Problem in rings of arbitrary positive characteristics. In contrast to fields, the characteristic of such a ring is not necessarily prime (e.g., the polynomial ring $\mathbb{Z}_{/6}[X]$ has characteristic $6$). Thus a difficulty is to describe the interaction of different prime divisors of the characteristic.
@@
Theorem 1.1 for rings of prime characteristics can be deduced from the case of fields following Derksen’s result. By [derksen2007skolem, Section 9], given any linear recurrence sequence $\gamma$ over a finitely generated ring of prime characteristic $p$, its zero set $\mathfrak{Z}(\gamma)\coloneqq\{n\in\mathbb{N}\mid\gamma_{n}=0\}$ is effectively $p$-normal (see Section 2 for a formal definition). For example, the linear recurrence sequence $\gamma$ over the polynomial ring $\mathbb{F}_{p}[X]$, defined by $\gamma_{n}=(X+1)^{n}-X^{n}-1$, admits the zero set $\{n=p^{k}\mid k\in\mathbb{N}\}$ [derksen2007skolem, Example 1.3].
@@
The formal definition of $p$-normal sets will be given in Section 2 (Definition 2.1). An example of $p$-normal sets is $\{1+5\cdot p^{2a}+p^{2b}\mid a,b\in\mathbb{N}\}$. In particular, these are special forms of $p$-automatic sets [wolper2000construction], and is a generalization of linear combinations of powers of $p$. Our key argument to proving Proposition 1.2 is a deep result of Dong and Shafrir [dong2025s, Theorem 1.3], who showed that the solution set to an S-unit equation over a $p^{e}$-torsion module is effectively $p$-normal.
The second obstacle is to compute the intersection $\mathfrak{Z}(\alpha_{1})\cap\cdots\cap\mathfrak{Z}(\alpha_{k})$ of the zero sets. Due to Proposition 1.2, this boils down to computing the intersection of $p_{i}$-normal sets for different primes $p_{1},\ldots,p_{k}$. In general, it is unknown whether one can effectively compute the intersection of $p_{i}$-automatic sets for different $p_{i}$’s [Hieronymi2022, albayrak2023quantitative]. However, we will show that this is possible in the special case of $p_{i}$-normal sets. In particular, we will prove that every intersection of $p_{i}$-normal sets can actually be written as a union of $p_{i}$-normal sets:
@@
Our main tool for proving Proposition 1.3 comes from a recent paper by Karimov, Luca, Nieuwveld, Ouaknine, and Worrell [karimov2025decidability], which gave an effective procedure to decide the existential fragment of Presburger arithmetic extended with two power predicates. For example, this allows an effective resolution of equations of the form $1+5\cdot p_{1}^{2a}+p_{1}^{2b}=3+p_{2}^{c}$, over the variables $a,b,c\in\mathbb{N}$. However, the effective procedure in [karimov2025decidability] does not extend to more than two power predicates. Therefore, additional insights are needed to compute the intersection of $p_{i}$-normal sets for $k\geq 3$ distinct primes $p_{1},\ldots,p_{k}$.
@@ 2.1 $p$-normal sets
We define $p$-normal subsets of $\mathbb{Z}^{n}$, $\mathbb{Z}$, and $\mathbb{N}$, following the convention established in [derksen2007skolem] and [derksen2015linear].
Definition 2.1 (reformulation of [derksen2015linear]).
@@
Lemma 2.4 ([derksen2007skolem, Lemma 9.5]).
@@
Lemma 2.5 (Special case of [derksen2015linear, Proposition 2.1]).
@@
All rings considered in this paper are commutative and unitary. We recall some standard definitions from commutative algebra [eisenbud2013commutative].
@@
    Let $I$ be an ideal in $R$. Then $I$ can always be written as a finite union of primary ideals: $I=P_{1}\cap\cdots\cap P_{n}$. Such a finite union is called a primary decomposition of $I$. Given a finite set of generators of $I$, a primary decomposition of $I$ can be effectively computed [rutman1992grobner].
@@
    Let $S\subset R$ be a set of non zero-divisors of $R$ and denote by $\widetilde{S}\coloneqq\{s_{1}s_{2}\cdots s_{k}\mid k\in\mathbb{N},s_{1},\ldots,s_{k}\in S\}$ its multiplicative closure. The localization $S^{-1}R$ is defined as the ring $\left\{\frac{r}{s}\;\middle|\;r\in R,s\in\widetilde{S}\right\}$, endowed with the operations $\frac{r_{1}}{s_{1}}+\frac{r_{2}}{s_{2}}=\frac{r_{1}s_{2}+r_{2}s_{1}}{s_{1}s_{2}}$ and $\frac{r_{1}}{s_{1}}\cdot\frac{r_{2}}{s_{2}}=\frac{r_{1}r_{2}}{s_{1}s_{2}}$. In particular, $S^{-1}R$ contains $R$ as a subring $\left\{\frac{r}{1}\;\middle|\;r\in R\right\}$. For example, the localization $\{2\}^{-1}\mathbb{Z}$ is the ring of rational numbers whose denominators are powers of $2$. Note that for the canonical map $R\rightarrow S^{-1}R$ to be injective, the set $S$ must not contain zero-divisors.
@@ 3 Skolem Problem in rings of positive characteristic
(c.f. [matsumura1989commutative, Theorem 1.3]). For each $i=1,\ldots,k$, consider the quotient $\mathcal{A}_{i}\coloneqq\mathcal{R}/p_{i}^{e_{i}}\mathcal{R}$ and the projection map $\rho_{i}\colon\mathcal{R}\rightarrow\mathcal{R}/p_{i}^{e_{i}}\mathcal{R}$. Then $\alpha_{i}\coloneqq\rho_{i}(\gamma)$ is a linear recurrence sequence over the ring $\mathcal{A}_{i}$. By Equation (3.1), we have $r=0$ if and only if $\rho_{i}(r)=0$ for all $i=1,\ldots,k$. Therefore
@@ 3.1 Effective $p_{i}$-normality of $\mathfrak{Z}(\alpha_{i})$
From now on we suppose that $f$ splits as a product $\prod_{i=1}^{t}(X-r_{i})^{d_{i}}$ over $\mathcal{A}$. When $\mathcal{A}$ is a field or an integral domain, it is a standard result [ouaknine2015linear] that the $n$-th term of the sequence $\alpha$ can be written as an exponential polynomial $\alpha_{n}=\sum_{i=1}^{t}r_{i}^{n}g_{i}(n)$ with $g_{i}(X)\in\mathcal{A}[X]$, (for example by considering the Jordan normal form of the transition matrix of $\alpha$). However this no longer holds in the presence of zero-divisors in the ring $\mathcal{A}$, since these zero-divisors can neither be inverted nor localized. To overcome this, we will use primary decomposition to reduce to the case where all zero-divisors are nilpotent.
@@
Since the intersection of $p$-normal subsets is effectively $p$-normal (Lemma 2.4), to show $\mathfrak{Z}(\alpha)$ is $p$-normal it suffices to show that each $\mathfrak{Z}(\tau_{j}(\alpha)),j=1,\ldots,s$ is $p$-normal. Thus, we are reduced to considering the case where $\mathcal{A}$ is of the form $A/P$, where $A$ is a ring with $p^{e}A=0$, and $P$ is a primary ideal of $A$. In particular, every zero-divisor of $\mathcal{A}=A/P$ is nilpotent (Definition 2.7(iv)).
@@
Suppose every zero-divisor of $\mathcal{A}$ is nilpotent, and the characteristic polynomial $f$ of $(\alpha_{n})_{n\in\mathbb{N}}$ splits as a product $\prod_{i=1}^{t}(X-r_{i})^{d_{i}}$ with $t\in\mathbb{N},r_{1},\ldots,r_{t}\in\mathcal{A}$. Then there is a finite, effectively computable set of non zero-divisors $S\subseteq\mathcal{A}$, numbers $u,N\in\mathbb{N}$, a subset $\mathcal{I}\subseteq\{1,\ldots,t\}$, and $c_{ij}\in S^{-1}\mathcal{A},i\in\mathcal{I};j=1,\ldots,u$, such that
$\alpha_{n}=\sum_{i\in\mathcal{I}}r_{i}^{n}\sum_{j=0}^{u}c_{ij}n^{j}$ (3.2)
@@
$\phi(Y)\coloneqq Y^{d}\cdot f\left(\frac{1}{Y}\right)=\prod_{i=1}^{t}(1-r_{i}Y)^{d_{i}}=1-a_{1}Y-\cdots-a_{d-1}Y^{d-1}-a_{d}Y^{d}.$
@@
$\sum_{i=1}^{t}\sum_{j=1}^{u}a_{ij}\cdot\frac{1}{(1-r_{i}Y)^{j}}$ (3.3)
for some $u\in\mathbb{N}$, $a_{ij}\in\widetilde{\mathcal{A}}[Y],1\leq i\leq t,0\leq j\leq u$. The proof is done in two steps.
@@
$\frac{1}{(1-r_{1}Y)(1-r_{2}Y)}$ $=\frac{1}{(1-r_{2}Y)^{2}}\cdot\frac{1}{\frac{1-r_{1}Y}{1-r_{2}Y}}$
$=\frac{1}{(1-r_{2}Y)^{2}}\cdot\frac{1}{1-\frac{(r_{1}-r_{2})Y}{1-r_{2}Y}}$
$=\frac{1}{(1-r_{2}Y)^{2}}\cdot\left(1+\frac{(r_{1}-r_{2})Y}{1-r_{2}Y}+\cdots+\frac{(r_{1}-r_{2})^{\ell-1}Y^{\ell-1}}{(1-r_{2}Y)^{\ell-1}}\right)$
$=\frac{1}{(1-r_{2}Y)^{2}}+\frac{(r_{1}-r_{2})Y}{(1-r_{2}Y)^{3}}+\cdots+\frac{(r_{1}-r_{2})^{\ell-1}Y^{\ell-1}}{(1-r_{2}Y)^{\ell+1}}.$
Divide both sides by $(1-r_{1}Y)^{d_{1}-1}(1-r_{2}Y)^{d_{2}-1}\prod_{i=3}^{t}(1-r_{i}Y)^{d_{i}}$, we obtain that $\frac{1}{\phi(Y)}=\frac{1}{\prod_{i=1}^{t}(1-r_{i}Y)^{d_{i}}}$ is equal to a sum of $\ell$ fractions
@@
$\frac{1}{\phi(Y)}=\sum_{i=1}^{t}\sum_{j=1}^{u}a_{ij}\cdot\frac{1}{(1-r_{i}Y)^{j}}.$
Consider a term $a_{ij}\cdot\frac{1}{(1-r_{i}Y)^{j}}$, we have
@@
$a_{ij}\cdot\frac{1}{(1-r_{i}Y)^{j}}=\sum_{n=0}^{\infty}\sum_{s=0}^{D_{ij}}\binom{j+n-s-1}{j-1}(b_{s}r_{i}^{-s})r_{i}^{n}Y^{n}$
and its $n$-th coefficient is some exponential-polynomial sum $r_{i}^{n}\sum_{j=0}^{u}c_{ij}n^{j}$. If $r_{i}$ is a zero-divisor and hence nilpotent, then $r_{i}^{\ell_{i}}=0$ for some effectively computable $\ell_{i}\in\mathbb{N}$, so $a_{ij}\cdot\frac{1}{(1-r_{i}Y)^{j}}$ is a polynomial of degree at most $\ell_{i}+D_{ij}$. This allows us to conclude the proof of the lemma by taking $N\coloneqq 1+\max_{i\not\in\mathcal{I},0\leq j\leq u}\{\ell_{i}+D_{ij}\}$. ∎
@@
$\alpha_{n}=\sum_{i\in\mathcal{I}}r_{i}^{n}\sum_{j=0}^{u}c_{ij}n^{j}$
for all $n\geq N$, with $c_{ij},r_{i}\in\mathcal{A}$ for all $i,j$. In particular, each $r_{i}$ is invertible.
A linear recurrence $\alpha$ is called simple if $u=0$ in the above expression. Since $p^{e}=0$ in $\mathcal{A}$, we have $n^{j}=(n+p^{e})^{j}$ in $\mathcal{A}$. Therefore the linear recurrence sequence $\alpha$ can be written as a union of $p^{e}$ different simple recurrences $\left(\alpha_{p^{e}n}\right)_{n\in\mathbb{N}},\left(\alpha_{p^{e}n+1}\right)_{n\in\mathbb{N}},\ldots,\left(\alpha_{p^{e}n+(p^{e}-1)}\right)_{n\in\mathbb{N}}$. In particular, we have
$\alpha_{p^{e}n+q}=\sum_{i\in\mathcal{I}}r_{i}^{p^{e}n+q}\sum_{j=0}^{u}c_{ij}(p^{e}n+q)^{j}=\sum_{i\in\mathcal{I}}\left(r_{i}^{p^{e}}\right)^{n}d_{i,q},$
with $d_{i,q}\coloneqq r_{i}^{q}\sum_{j=0}^{u}c_{ij}q^{j}$. Since $p$-normal sets are closed under finite union and affine transformations, it suffices to show that each simple recurrence $\left(\alpha_{p^{e}n+q}\right)_{n\in\mathbb{N}}$ has an effectively $p$-normal zero set.
Therefore we now without loss of generality replace each $r_{i}$ with $r_{i}^{p^{e}}$ and consider each simple recurrence. Let
@@
Write $\mathcal{I}$ as $\{1,\ldots,t\}$ for some $t\in\mathbb{N}$. We can give $\mathcal{A}$ a $\mathbb{Z}[X_{1}^{\pm},\ldots,X_{t}^{\pm}]$-module structure by letting $X_{i}$ act as $r_{i}$. Denote by $\mathcal{M}$ the submodule generated by $c_{1},\ldots,c_{t}$, we have $p^{e}\mathcal{M}=0$. Using the standard algorithm in commutative algebra (see [baumslag1981computable, Theorem 2.14] or [baumslag1994algorithmic, Theorem 2.6]) we can compute a finite presentation of $\mathcal{M}$ as a $\mathbb{Z}[X_{1}^{\pm},\ldots,X_{t}^{\pm}]$-module. Then, the equation $\sum_{i=1}^{t}c_{i}r_{i}^{z}=0$ in $\mathcal{A}$ is equivalent to the equation $\sum_{i=1}^{t}X_{i}^{z}\cdot c_{i}=0$ in the $\mathbb{Z}[X_{1}^{\pm},\ldots,X_{t}^{\pm}]$-module $\mathcal{M}$.
@@
The previous discussion showed that we can reduce to the case where $\alpha_{n}=\sum_{i=1}^{t}c_{i}r_{i}^{n}$ for all $n\geq N$, where each $r_{i}$ is invertible. For each $0\leq n<N$, we can verify whether $\alpha_{n}=0$ individually. Lemma 3.3 showed that the zero set

### v3

@@ 1 Introduction and Main Results
The Skolem Problem in a commutative ring $\mathcal{R}$ is the problem of determining whether a given linear recurrence sequence $(\gamma_{n})_{n\in\mathbb{N}}$ over $\mathcal{R}$ contains a zero term. The Skolem Problem is deeply related to central questions in program verification [OW15, AKK+21], control theory [BT00, FOP+19], dynamical systems [KL86, AAG+15], and number theory [BEL06, LMN+23]. Despite its simple formulation, decidability of the Skolem Problem over $\mathbb{Z}$ (and equivalently, $\mathbb{Q}$) is a major open problem in mathematics and computer science: see [MAH35, LEC53, TMS84, BP02, LLN+22, LOW25] for a selected list of partial results and recent progress. Indeed, the celebrated Skolem-Mahler-Lech theorem [SKO34] shows that the zero set of a linear recurrence sequence over $\mathbb{Z}$ is a union of a finite set and finitely many arithmetic progressions. However, the Skolem-Mahler-Lech theorem is not effective, meaning there is currently no known computable bound on this finite set, and hence the decidability of the Skolem Problem over the ring $\mathbb{Z}$ remains open.
In contrast, decidability of the Skolem Problem in fields of prime characteristic (such as $\mathbb{F}_{p}(X)$) was shown by Derksen [DER07], who gave an effective description of the zero set as a $p$-normal set. In this paper we continue Derksen’s work in this direction by showing decidability of the Skolem Problem in rings of arbitrary positive characteristics. In contrast to fields, the characteristic of such a ring is not necessarily prime (e.g., the polynomial ring $\mathbb{Z}_{/6}[X]$ has characteristic $6$). Thus a difficulty is to describe the interaction of different prime divisors of the characteristic.
@@
Theorem 1.1 for rings of prime characteristics can be deduced from the case of fields following Derksen’s result. By [DER07, Section 9], given any linear recurrence sequence $\gamma$ over a finitely generated ring of prime characteristic $p$, its zero set $\mathfrak{Z}(\gamma)\coloneqq\{n\in\mathbb{N}\mid\gamma_{n}=0\}$ is effectively $p$-normal (see Section 2 for a formal definition). For example, the linear recurrence sequence $\gamma$ over the polynomial ring $\mathbb{F}_{p}[X]$, defined by $\gamma_{n}=(X+1)^{n}-X^{n}-1$, admits the zero set $\{n=p^{k}\mid k\in\mathbb{N}\}$ [DER07, Example 1.3].
@@
The formal definition of $p$-normal sets will be given in Section 2 (Definition 2.1). An example of $p$-normal sets is $\{1+5\cdot p^{2a}+p^{2b}\mid a,b\in\mathbb{N}\}$. In particular, these are special forms of $p$-automatic sets [WB00], and is a generalization of linear combinations of powers of $p$. Our key argument to proving Proposition 1.2 is a deep result of Dong and Shafrir [DS26, Theorem 1.3], who showed that the solution set to an S-unit equation over a $p^{e}$-torsion module is effectively $p$-normal.
The second obstacle is to compute the intersection $\mathfrak{Z}(\alpha_{1})\cap\cdots\cap\mathfrak{Z}(\alpha_{k})$ of the zero sets. Due to Proposition 1.2, this boils down to computing the intersection of $p_{i}$-normal sets for different primes $p_{1},\ldots,p_{k}$. In general, it is unknown whether one can effectively compute the intersection of $p_{i}$-automatic sets for different $p_{i}$’s [HS22, AB23]. However, we will show that this is possible in the special case of $p_{i}$-normal sets. In particular, we will prove that every intersection of $p_{i}$-normal sets can actually be written as a union of $p_{i}$-normal sets:
@@
Our main tool for proving Proposition 1.3 comes from a recent paper by Karimov, Luca, Nieuwveld, Ouaknine, and Worrell [KLN+25], which gave an effective procedure to decide the existential fragment of Presburger arithmetic extended with two power predicates. For example, this allows an effective resolution of equations of the form $1+5\cdot p_{1}^{2a}+p_{1}^{2b}=3+p_{2}^{c}$, over the variables $a,b,c\in\mathbb{N}$. However, the effective procedure in [KLN+25] does not extend to more than two power predicates. Therefore, additional insights are needed to compute the intersection of $p_{i}$-normal sets for $k\geq 3$ distinct primes $p_{1},\ldots,p_{k}$.
@@ 2.1 $p$-Normal Sets
We define $p$-normal subsets of $\mathbb{Z}^{n}$, $\mathbb{Z}$, and $\mathbb{N}$, following the convention established in [DER07] and [DM15].
Definition 2.1 (reformulation of [DM15]).
@@
Lemma 2.4 ([DER07, Lemma 9.5]).
@@
Lemma 2.5 (Special case of [DM15, Proposition 2.1]).
@@ 2.2 Commutative Algebra
All rings considered in this paper are commutative and unitary. We recall some standard definitions from commutative algebra [EIS13].
@@
    Let $I$ be an ideal in $R$. Then $I$ can always be written as a finite union of primary ideals: $I=P_{1}\cap\cdots\cap P_{n}$. Such a finite union is called a primary decomposition of $I$. Given a finite set of generators of $I$, a primary decomposition of $I$ can be effectively computed [RUT92].
@@
    Let $S\subset R$ be a set of non zero-divisors of $R$ and denote by $\widetilde{S}\coloneqq\{s_{1}s_{2}\cdots s_{k}\mid k\in\mathbb{N},s_{1},\ldots,s_{k}\in S\}$ its multiplicative closure. The localization $S^{-1}R$ is defined as the ring $\left\{\frac{r}{s}\;\middle|\;r\in R,s\in\widetilde{S}\right\}$, with $\frac{r_{1}}{s_{1}}=\frac{r_{2}}{s_{2}}$ if and only if $s_{1}r_{2}-s_{2}r_{1}$ is a zero-divisor. This ring is endowed with the operations $\frac{r_{1}}{s_{1}}+\frac{r_{2}}{s_{2}}=\frac{r_{1}s_{2}+r_{2}s_{1}}{s_{1}s_{2}}$ and $\frac{r_{1}}{s_{1}}\cdot\frac{r_{2}}{s_{2}}=\frac{r_{1}r_{2}}{s_{1}s_{2}}$. In particular, $S^{-1}R$ contains $R$ as a subring $\left\{\frac{r}{1}\;\middle|\;r\in R\right\}$. For example, the localization $\{2\}^{-1}\mathbb{Z}$ is the ring of rational numbers whose denominators are powers of $2$. Note that in order for the canonical map $R\rightarrow S^{-1}R$ to be injective, the set $S$ must not contain zero-divisors.
@@ 3 Skolem Problem in Rings of Positive Characteristic
(c.f. [MAT89, Theorem 1.3]). For each $i=1,\ldots,k$, consider the quotient $\mathcal{A}_{i}\coloneqq\mathcal{R}/p_{i}^{e_{i}}\mathcal{R}$ and the projection map $\rho_{i}\colon\mathcal{R}\rightarrow\mathcal{R}/p_{i}^{e_{i}}\mathcal{R}$. Then $\alpha_{i}\coloneqq\rho_{i}(\gamma)$ is a linear recurrence sequence over the ring $\mathcal{A}_{i}$. By Equation (3.1), we have $r=0$ if and only if $\rho_{i}(r)=0$ for all $i=1,\ldots,k$. Therefore
@@ 3.1 Effective $p_{i}$-Normality of $\mathfrak{Z}(\alpha_{i})$
From now on we suppose that $f$ splits as a product $\prod_{i=1}^{t}(X-r_{i})^{d_{i}}$ over $\mathcal{A}$. When $\mathcal{A}$ is a field or an integral domain, it is a standard result [OW15] that the $n$-th term of the sequence $\alpha$ can be written as an exponential polynomial $\alpha_{n}=\sum_{i=1}^{t}r_{i}^{n}g_{i}(n)$ with $g_{i}(X)\in\mathcal{A}[X]$, (for example by considering the Jordan normal form of the transition matrix of $\alpha$). However this no longer holds in the presence of zero-divisors in the ring $\mathcal{A}$, since these zero-divisors can neither be inverted nor localized. To overcome this, we will use primary decomposition to reduce to the case where all zero-divisors are nilpotent.
@@
Since the intersection of $p$-normal subsets is effectively $p$-normal (Lemma 2.4), to show $\mathfrak{Z}(\alpha)$ is $p$-normal it suffices to show that each $\mathfrak{Z}(\tau_{j}(\alpha)),j=1,\ldots,s$ is $p$-normal. Thus, we are reduced to considering the case where $\mathcal{A}$ is of the form $A/P$, where $A$ is a ring with $p^{e}A=0$, and $P$ is a primary ideal of $A$. In particular, every zero-divisor of $\mathcal{A}=A/P$ is nilpotent (Definition 2.7(iv)). For $n\in\mathbb{Z},k\in\mathbb{N}$, let $\binom{n}{k}$ denote the binomial coefficient.
@@
Suppose every zero-divisor of $\mathcal{A}$ is nilpotent, and the characteristic polynomial $f$ of $(\alpha_{n})_{n\in\mathbb{N}}$ splits as a product $\prod_{i=1}^{t}(X-r_{i})^{d_{i}}$ with $t\in\mathbb{N},r_{1},\ldots,r_{t}\in\mathcal{A}$. Then there is a finite, effectively computable set of non zero-divisors $S\subseteq\mathcal{A}$, numbers $u,N\in\mathbb{N}$, a subset $\mathcal{I}\subseteq\{1,\ldots,t\}$, and $c_{ik}\in S^{-1}\mathcal{A},i\in\mathcal{I};k=1,\ldots,u$, such that
$\alpha_{n}=\sum_{i\in\mathcal{I}}r_{i}^{n}\sum_{k=0}^{u}\binom{n}{k}c_{ik}$ (3.2)
@@
$\phi(Y)\coloneqq Y^{d}f\left(\frac{1}{Y}\right)=\prod_{i=1}^{t}(1-r_{i}Y)^{d_{i}}=1-a_{1}Y-\cdots-a_{d}Y^{d}.$
@@
$\sum_{i=1}^{t}\sum_{j=1}^{v}a_{ij}\cdot\frac{1}{(1-r_{i}Y)^{j}}$ (3.3)
for some $v\in\mathbb{N}$, $a_{ij}\in\widetilde{\mathcal{A}}[Y],1\leq i\leq t,0\leq j\leq v$. The proof is done in two steps.
@@
$\frac{1}{1-r_{1}Y}$ $=\frac{1}{1-r_{2}Y}\cdot\frac{1}{\frac{1-r_{1}Y}{1-r_{2}Y}}$
$=\frac{1}{1-r_{2}Y}\cdot\frac{1}{1-\frac{(r_{1}-r_{2})Y}{1-r_{2}Y}}$
$=\frac{1}{1-r_{2}Y}\cdot\left(1+\frac{(r_{1}-r_{2})Y}{1-r_{2}Y}+\cdots+\frac{(r_{1}-r_{2})^{\ell-1}Y^{\ell-1}}{(1-r_{2}Y)^{\ell-1}}\right)$
$=\frac{1}{1-r_{2}Y}+\frac{(r_{1}-r_{2})Y}{(1-r_{2}Y)^{2}}+\cdots+\frac{(r_{1}-r_{2})^{\ell-1}Y^{\ell-1}}{(1-r_{2}Y)^{\ell}}.$
Divide both sides by $(1-r_{1}Y)^{d_{1}-1}(1-r_{2}Y)^{d_{2}}\prod_{i=3}^{t}(1-r_{i}Y)^{d_{i}}$, we obtain that $\frac{1}{\phi(Y)}=\frac{1}{\prod_{i=1}^{t}(1-r_{i}Y)^{d_{i}}}$ is equal to a sum of $\ell$ fractions
@@
$\frac{1}{\phi(Y)}=\sum_{i=1}^{t}\sum_{j=1}^{v}a_{ij}\cdot\frac{1}{(1-r_{i}Y)^{j}}$
for some $v\in\mathbb{N}$. Consider a term $a_{ij}\cdot\frac{1}{(1-r_{i}Y)^{j}}$, we have
@@
$a_{ij}\cdot\frac{1}{(1-r_{i}Y)^{j}}=\sum_{n=0}^{\infty}\sum_{s=0}^{D_{ij}}\binom{j+n-s-1}{j-1}(b_{s}r_{i}^{-s})r_{i}^{n}Y^{n}.$
Its $n$-th coefficient can be written as some sum $r_{i}^{n}\sum_{k=0}^{u}\binom{n}{k}c_{ik}$, by applying the identity $\binom{n+z}{j-1}=\sum_{k=0}^{j-1}\binom{z}{j-1-k}\binom{n}{k}$. If $r_{i}$ is a zero-divisor and hence nilpotent, then $r_{i}^{\ell_{i}}=0$ for some effectively computable $\ell_{i}\in\mathbb{N}$, so $a_{ij}\cdot\frac{1}{(1-r_{i}Y)^{j}}$ is a polynomial of degree at most $\ell_{i}+D_{ij}$. This allows us to conclude the proof of the lemma by taking $N\coloneqq 1+\max_{i\not\in\mathcal{I},0\leq j\leq v}\{\ell_{i}+D_{ij}\}$. ∎
@@
$\alpha_{n}=\sum_{i\in\mathcal{I}}r_{i}^{n}\sum_{k=0}^{u}\binom{n}{k}c_{ik}$
for all $n\geq N$, with $c_{ik},r_{i}\in\mathcal{A}$ for all $i,k$. In particular, each $r_{i}$ is invertible.
A linear recurrence $\alpha$ is called simple if $u=0$ in the above expression. Since $p^{e}=0$ in $\mathcal{A}$, it is easy to see that for all $w\geq e+k!$ we have $\binom{n}{k}=\binom{n+p^{w}}{k}$ in $\mathcal{A}$. Let $w\coloneqq e+u!$, then the linear recurrence sequence $\alpha$ can be written as a union of $p^{w}$ different simple recurrences $\left(\alpha_{p^{w}n}\right)_{n\in\mathbb{N}},\left(\alpha_{p^{w}n+1}\right)_{n\in\mathbb{N}},\ldots,\left(\alpha_{p^{w}n+(p^{w}-1)}\right)_{n\in\mathbb{N}}$. In particular, we have
$\alpha_{p^{w}n+q}=\sum_{i\in\mathcal{I}}r_{i}^{p^{w}n+q}\sum_{k=0}^{u}\binom{p^{e}n+q}{k}c_{ik}=\sum_{i\in\mathcal{I}}\left(r_{i}^{p^{w}}\right)^{n}d_{i,q},$
with $d_{i,q}\coloneqq r_{i}^{q}\sum_{k=0}^{u}\binom{q}{k}c_{ik}$. Since $p$-normal sets are closed under finite union and affine transformations, it suffices to show that each simple recurrence $\left(\alpha_{p^{w}n+q}\right)_{n\in\mathbb{N}}$ has an effectively $p$-normal zero set.
Therefore we now without loss of generality replace each $r_{i}$ with $r_{i}^{p^{w}}$ and consider each simple recurrence. Let
@@
Write $\mathcal{I}$ as $\{1,\ldots,t\}$ for some $t\in\mathbb{N}$. We can give $\mathcal{A}$ a $\mathbb{Z}[X_{1}^{\pm},\ldots,X_{t}^{\pm}]$-module structure by letting $X_{i}$ act as $r_{i}$. Denote by $\mathcal{M}$ the submodule generated by $c_{1},\ldots,c_{t}$, we have $p^{e}\mathcal{M}=0$. Using the standard algorithm in commutative algebra (see [BCM81, Theorem 2.14] or [BCR94, Theorem 2.6]) we can compute a finite presentation of $\mathcal{M}$ as a $\mathbb{Z}[X_{1}^{\pm},\ldots,X_{t}^{\pm}]$-module. Then, the equation $\sum_{i=1}^{t}c_{i}r_{i}^{z}=0$ in $\mathcal{A}$ is equivalent to the equation $\sum_{i=1}^{t}X_{i}^{z}\cdot c_{i}=0$ in the $\mathbb{Z}[X_{1}^{\pm},\ldots,X_{t}^{\pm}]$-module $\mathcal{M}$.
@@
The previous discussion shows that we can reduce to the case where $\alpha_{n}=\sum_{i=1}^{t}c_{i}r_{i}^{n}$ for all $n\geq N$, where each $r_{i}$ is invertible. For each $0\leq n<N$, we can verify whether $\alpha_{n}=0$ individually. Lemma 3.3 showed that the zero set